OpenStudy (1645323):

Find the number of solutions y = 3x2 - 4x - 5 Select one: a. 2 b. 0 c. 1 d. 2

jimthompson5910 (jim_thompson5910):

hint: the equation ax^2 + bx + c = 0 will either have * two real solutions * one real solution * no real solutions To determine which scenario happens, we use the formula D = b^2 - 4ac, which is the discriminant formula. We then use these rules * If D > 0, then ax^2 + bx + c = 0 has 2 real solutions * If D = 0, then ax^2 + bx + c = 0 has 1 real solution * If D < 0, then ax^2 + bx + c = 0 has no real solutions